OpenStudy (anonymous):

what is the solution set of the inequality -5x + (-3) <2x -17

OpenStudy (lncognlto):

Well, what do you get when you solve for x?

OpenStudy (anonymous):

i don't kno

OpenStudy (lncognlto):

Ok, so -5x + (-3) <2x -17 = -3 + 17 < 2x + 5x = 14 < 7x = x > ?